OpenStudy (anonymous):

Katie has 18 fish in her tank. Six of the fish are orange and the rest are yellow. What is the ratio of yellow fish to orange fish? A. 1 : 2 B. 2 : 3 C. 2 : 1 D. 3 : 2

OpenStudy (anonymous):

do u think u can help me

OpenStudy (anonymous):

I think so let me just read over the problem

OpenStudy (anonymous):

okay, so you have 18 fish, and 6 of the fish are orange. 18-6=12, so there are 12 yellow fish. the ratio of yellow to orange would then be 12:6. however, that's not one of the answer choices. you can simplify a ratio the same way you simplify a fraction though. so the answer would then be 2:1

OpenStudy (anonymous):

\[\frac{18-6}{6}=2 \]